### v1

# A scanning apparatus to detect the spectral directivity of optically-emissive materials

A scanning apparatus to detect the spectral directivity of optically-emissive materials
††journal: opticajournal††articletype: Research ArticleAn apparatus that records the optical spectrum of emissive materials as a function of the polar coordinate angles is reported. The ability for the device to characterize the directivity of a light source over the optical spectrum is demonstrated. The angular dependence of an electrically-driven LED with a hemispherical diffuser cap was characterized. Optically pumped materials that emitted both fluorescence and amplified spontaneous emission (ASE) were also characterized.
1 Introduction
The passive linear optical properties of a material can be characterized by detecting changes to an incident probe beam using methods such as reflection, transmission, and absorption spectroscopy.[1] Active techniques involve probing the material upon exciting it with i) an external light source,[2] ii) electrically-driven charge carriers,[3] or iii) chemical reactions.[4] Measuring the directivity over a range of the light spectrum exiting a sample after relaxing from an excited state is the primary focus of this paper.
When an a low population density of excited states of molecular ensembles release electromagnetic energy, the phenomena is referred to as fluorescence.[5] Dispersing a low concentration of fluorophores in a transparent matrix is the simplest way to ensure a low population density of noninteracting molecules. Fluorescent molecules with long excited-state lifetimes doped in a homogeneous liquid will act as an unpolarized source with the same volumetric shape as the absorbed light. Changing the surface topology of a sample can affect the directivity (e.g., surface relief grating).[6, 7] Furthermore, refractive index-mismatched internal structures can create additional reflecting surfaces which dictate the directivity of light.[8]
When a high concentration of fluorophores is pumped by a strong light source, a population inversion density can be achieved. Many fluorescent molecules behave like a 4-level model.[9, 10] If the path length of an emitted photon through the gain region of excited molecules is long enough, then amplified spontaneous emission (ASE) can occur.[11] Traditionally, a long strip of material is excited to produce ASE.[12, 13] Aperiodic scattering surfaces inside the gain region can increase the effective path length of emitted photons which can create directed ASE.[14, 15, 16, 17] The narrow spectral profile for this type of enhanced ASE generation is similar to the spectrum produced by multimode laser emission from an “intensity feedback” or “non-resonant feedback” random laser.[18, 19, 20, 21] Natural and fabricated materials can possess a highly periodic internal structure.[22, 23, 24, 25, 26, 27] When the internal structure is highly periodic, laser light can be emitted (e.g., multilayer distributed feedback laser).[28, 29, 30, 31] At high thresholds, even aperiodic structures can produce laser emission, which has been referred to as a “resonant feedback” random laser.[32, 33, 34] The directivity from the complex structures and properties given in the latter examples can be probed by measuring the spectrum of the emitted light over the polar and azimuth angles. This paper presents a spectral scanner, which collects light via a miniature spectrometer while scanning through the polar and azimuth angles to characterize the emissive properties of simple and complex-structured materials and devices.
2 Experimental methods
2.1 Apparatus
Automated recordings of emission spectra scanned over the azimuth and polar angles were performed by an in-house built apparatus. An aluminum sample stage was at the center of the apparatus and temperature controlled by running liquid through a bored cavity. The sample stage was secured by a post that ran through a bore hole in a rotation stage that swept the azimuth angles as shown in Figure 1. A circular plate was attached to the rotation stage. An off-center rotation stage was attached to sweep the polar angles. The circular plate was used to mark rotation angles to initially configure the azimuth rotation stage, but a horizontal jib arm holding the second rotation stage was also found to work in the same capacity. A boom arm was attached to the second rotation stage. A lever arm was attached at the top of the boom arm to secure a sensor and its alignment components.
A tunable double lens tube was used to zoom and focus in on the sample stage. Adjustments to the lens tube’s orientation were made by mounting it on a kinematic mirror mount. The kinematic mirror mount is mounted on a slot and the distance between zoom lens tube and the sample plate can be adjusted, A camera was placed in the image location for alignment purposes. After alignment, the camera was replaced by a m-core multimode optical patch fiber. The fiber was connected to a portable spectrometer mounted to the boom arm. We initially attempted to secure one end of a long fiber to the apparatus and the other end to a spectrometer mounted on the optical table, but fiber deformations from apparatus rotations caused the relative intensity of the signal to change. By mounting the spectrometer to the boom arm with a rigid fiber patch cord, loss through the optical fiber was held constant. A slot was located between the lens tube and the detector position which allowed for 1" filters (chromatic, polarization, etc.).
The motors on the apparatus were controlled by a CNC shield attached to an Arduino Uno loaded with a GRBL library. The Arduino was connected to a computer via USB port. The spectrometer was also connected to the computer. A Python script recorded spectra with the Seabreeze package, and it also controlled the apparatus motors. A third motor was connected to the CNC shield which rotated a halfwave plate as shown in Figure 1. The pump beam’s polarization was rotated and sent through a polarizing beam cube splitter. Pulse energy was adjusted using this halfwave plate and polarizer combination. A relay was also connected by USB port and incorporated into the same Python code that controlled the entire apparatus. The relay controlled a shutter, where the shutter was open for data collection of materials excited by the external pump laser and closed when the apparatus rotated during the scanning procedure. The shutter reduced fluorophore photodegradation during a scan.
2.2 Alignment
The sample stage can be placed at the approximate center of the azimuth rotating stage bore hole using -positioning stage at the base of the apparatus. The sample stage height can be adjusted so the sample plate is placed approximately lower than center of the polar rotating stage bore hole. Later on, the center of the azimuth rotation stage on the sample plate can be found and the sample will be placed on the spot.
Before connecting the optical fiber, a camera was first positioned behind the adjustable lens tube. The camera was used to finely align the system via an iterative technique. First the lens system has to be aligned with the arm. Placing the camera at a polar angle of and rotating the azimuth plate, the center spot of the azimuth rotation stage can be found. We then marked the spot on the computer screen. We noticed that the sample plate may not be perfectly at the center of the azimuth rotation stage. Using -positioning stage at the base of the apparatus to move the sample plate and then marked the center of the rotation stage on the sample plate. This spot is the sample position which will be aligned with the center of the camera. We can adjust the camera position so the sample position is in the center of the camera. After that, the apparatus then swung around . If the sample position rotated in a small circle or didn’t move, the kinematic mount was adjusted until the sample remained in the same position through the entire scan.
After the sample position was aligned in -plane. A LED or a droplet (sample) can be placed in the center to continue next alignment step. The detector was placed at a polar angle of . Then we adjusted the sample plate height to place the top of the sample in the center of the camera. Then the apparatus swung around and if the sample didn’t move or move in a line, then we can use the kinematic mount. A final check was performed by scanning the entire set of azimuth and polar angles to confirm the sample position remained constant from the camera perspective. The camera was then replaced by the fiber feeding into the spectrometer attached to the boom arm.
2.3 Experiment
The apparatus measures the directivity of an emissive material/device over the spectral range recorded by the fiber spectrometer. Two types of samples were recorded, 1) an electrically-driven light emitting diode (LED) with a diffuser cap and 2) a fluorescein-doped water droplet. The electrical current for the LED and the pump beam pulse energy for the fluorescein-doped water were held constant for the duration of their scans. All scans were stepped in the azimuth from to . They scans were stepped along the co-latitude from to .
The two different sizes of droplet samples were measured under the same geometric condition. For the first one, a large dye-doped water droplet was placed on a hydrophobic parafilm surface with a pump beam illuminating the sample from above. A second droplet was approximately 4 times smaller than the first droplet and tested by the similar size pump beam. A nm pulsed laser beam with a ns duration was used for this experiment.
3 Results
3.1 Diffuser-capped LED
Our system was aligned through the process described in Section 2.2. In an attempt to confirm the alignment, an LED with a diffuser cap was placed at the center of the sample stage. An ideal spherical diffuser cap would mimic a spherically symmetric light source, which would confirm the alignment; however, the spherical diffuser capping the LED did not have spherical light-emissive symmetry.
At high polar angles, a hot spot was formed along the azimuth as depicted in the intensity heat map shown in Figure 2(a). The nonuniform directivity could have been caused by an alignment issue, so the sample was rotated by and a new scan was performed. As shown in Figure 2(b), the recorded hot spot rotated in the azimuthal direction by the same angle that the sample was rotated. This azimuth rotation of the sample confirmed the hot spot was indeed sourced from the diffuser-capped LED as opposed to an issue with the system alignment.
As expected, there was no distinguishable difference between the peak wavelength and spectral width of the LED’s emitted light. The amplitude of the spectra shown in Figure 2(c-f) vary in amplitude only. There is a significant trend for the intensity’s dependence on the polar angle as shown in Figure 2(c-d). The detected signal is largest at near the top of the LED and significantly weakens as the detector collects light near the horizon. A hot spot is clearly identifiable as a function of the azimuthal angle. The peak amplitude of the spectra, shown in Figure 2(e-f), shifted by the same angle as the rotated sample. The LED has no significant process for emitting directed energy, and the diffuser cap is fabricated from a passive achromatic material. Therefore, as expected for a linear optical source, the spectral features remain constant while only the amplitude of the emitted light depends on the spherical angles. This result implies that no directed energy occurs at the source or during the passive diffusion through the cap, as expected.
3.2 Large water droplet: center pumped
The diffuser-capped, electrically-pumped, LED light source in the previous section had hot spots from aspherical light emission. A wt.% fluorescein/water solution was used in an attempt to create a better approximation to a spherical-emitting light source. The droplet was placed on parafilm, where the solution/air interface above the parafilm surface formed an approximate hemisphere. The beam spot was in diameter, which illuminated the majority of the liquid hemisphere’s surface. The liquid cooled stage was measured to be C while the lab was at a typical laboratory room temperature and humidity. The cooled surface resulted in a near steady state droplet size due to the competition between evaporation and condensation.
There is an increased emission intensity at higher co-latitude (lower latitude) for the droplet. This trend in emission as a function of polar angle is opposite of the diffuser-capped LED, where the spectra shown in Figure 3(a) can be visually compared those given in Figure 2(c).
The noise associated with the spectrum is also greater in the droplet at low emission intensity levels as compared to the diffuser-capped LED. The azimuthal dependence shown in Figure 2(b) also has some shot-to-shot variations throughout the scan. Some small variations in the beam profile could result in an azimuth-dependent emission; however, shot-to-shot noise from the pump beam also contribute to the variations in the azimuthal dependence of the emission intensity. The shot-to-shot noise of an Nd:YAG’s fundamental frequency is reasonable low. The shot-to-shot uncertainty of the pump is increased when a portion of the nm pump beam’s photons are frequency doubled by a beta Barium Borate (BBO) crystal followed by sum-frequency generation of the nm and nm to yield a nm beam. This nonlinear, macroscopic, cascading process results in larger variations between pulse energies for the nm beam entering an optical parametric oscillator (OPO). For the OPO stage, the residual nm and nm light is attenuated by a set of dichroic mirrors. The remaining nm light is injected into a cavity via another dichroic mirror with a long BBO crystal between two mirrors. The amplified output of the OPO for wavelength selection at nm also amplifies the shot-to-shot uncertainty of the pulse energy. Thus, the uncertainty between each shot appears to be relatively large.
Averaging out the uncertainty between recorded shots taken while the azimuth is scanned, a slow change of intensity is observed through the scan as shown in Figure 3(b). The output intensity has some preferred azimuthal emission direction. If the preferred direction were biaxial, then the cause would undoubtedly be from the pump beam’s polarization. The excited state lifetime of fluorescein is relatively large to the pump pulse duration, making a polarization dependent emission unlikely. Indeed, we found that the small, preferred, azimuth emission was unidirectional, which was caused by a uncertainties in the profile and center-alignment of the OPO pump beam. Although the overall signal is relatively small compared to the single-shot noise of the spectrometer, there is a clear direction of preferred emission. The ASE clearly dominates the fluorescence signal. The small fluorescence peak is observed at a nm wavelength while the peak of the red-shifted ASE profile is much greater at nm.
3.3 Small water droplet: off-center pumped
The directivity of the emission spectra for a smaller droplet of wt.% fluorescein/water resting on parafilm was also scanned. The droplet diameter was 2 times smaller than droplet in Section 3.2 which leads to 4 times smaller in size. The pump beam size was approximately same as the large droplet which is m. The pump pulse energy and compact gain region resulted in an above threshold mixture of ASE and fluorescence. Similar to the large droplet described in Section 3.2, the emission intensity increased with the co-latitude as shown in Figure 4(a).
The focused beam was slightly off center in this case which caused an asymmetric emission profile as measured in the azimuthal direction. Figure 4(b) shows emission spectra as a function of the azimuth angle for a fixed co-latitude. The small deviation from the vertical axis centered on the hemispherical droplet had significant effects on the azimuthal dependence of the emission intensity and type of radiative process. The fluorescence and ASE emission intensities were comparable when viewed from an azimuth angle of . As the detector scanned around the sample, the ASE began to dominate the emitted light signal. As with the large droplet, there was a significant amount of uncertainty caused by the shot-to-shot deviations of the OPO pulse energy. Even after considering this uncertainty, there is a clear trend associated with the emission intensity and the azimuthal angle.
4 Discussion
The diffuser-capped LED was initially intended to confirm alignment based on an assumed spherical symmetry for light emission from cosine correction across the hemispherical diffuser cap. In hindsight, the polar angle dependence can be easily explained when assuming a highly directional LED light source directed towards the top of the diffuser cap from the inside. The hot spot identified in the heat map was shown to be caused by the source as opposed to an error in alignment of the apparatus. Additional rotations beyond those shown in Figures 2(a) and 2(b) further confirmed the LED source had an azimuthal dependence on the emission intensity. Thus, the apparatus can be used to characterize electrically-driven light sources.
The high dye concentration in the water droplets was chosen because of the sensitivity of ASE emission. The ASE process occurs when 1) a population inversion density is present and 2) the path length of an emitted beam is long enough to amplify the emitted light. The amplification of the emission when the illumination was nearly cylindrically symmetric resulted in a near azimuthal symmetry for the emission; however, the uncertainty caused by shot-to-shot intensity variations of the OPO pump introduced some noise-like variance between each recorded spectrum which can be observed in Figure 3(a). The directivity was found to be independent of pump polarization.
The characterization for the off-center pumped droplet provided interesting insights. One might expect the ASE and fluorescence intensities to be weighted evenly with respect to each other at all emission angles. Instead we observed different directivities associated with the two emission processes. The results from this last study also illustrate the importance of pump beam characteristics in addition to apparatus alignment when scanning optically pumped light sources.
Although the apparatus was initially fabricated for use with a miniature spectrometer, other detectors could take its place. For example, a fast photodiode could be mounted to collect transient data, where emitted pulses differ between fluorescence, ASE, and lasing. This was not attempted with the current setup, where there was already a large uncertainty in the OPO’s shot-to-shot pulse energy appearing with spectrometer data. It may be of interest to test other sources such as a rhodamine derivative solution pumped by a frequency-doubled Nd:YAG, which should result in lower pulse energy deviations.
5 Conclusion
This paper describes an apparatus that scans the spectrum of emitted light as a function of the spherical coordinate angles. Alignment of the apparatus was achieved through the use of a camera which imaged the light source as it swept through a series of angles. The apparatus was found to be well-suited for characterization of LED light sources as well as emissive materials that require optical pumping. Beyond the example materials characterized in Section 3, the directivity of many types of emissive materials can be characterized. Additional examples include random lasers, light sources surrounded by dispersive elements, and complex biological materials stained with fluorescent molecules.
6 Back matter
Funding This material is based upon work supported by the National Science Foundation, Directorate for Mathematical and Physical Sciences (Grant No. 2337595).
Acknowledgment The authors thank Mike Kalman for his role as machinist during the build of the apparatus described in this paper.
Disclosures The presented device is granted with a USA Patent 12203850 on Jan 21 2025.
Data availability Data underlying the results presented in this paper are not publicly available at this time but may be obtained from the authors upon reasonable request.